# Adaptive Federated Learning with Auto-Tuned Clients

### v1

@@ 4.1 Results
Table 1: Experimental results based on the settings detailed in Section 4. Performance difference within 0.5% of the best result for each task are shown in bold. Subscripts${}_{\downarrow(x.x)}$ is the performance difference from the best result and is highlighted in pink when the difference is bigger than 2%. The symbol ($\downarrow$) appended to SGD and SGDM indicates step-wise learning rate decay, where the step sizes are divided by $10$ after 50%, and another by $10$ after 75% of the total rounds.
@@
Three illustrative cases are visualized in Figure 2. We remind that the hyperparameters are tuned for the case of CIFAR-10 trained with ResNet-18, with $\alpha=0.1$ (Figure 2(A)). For this task, with $\alpha=1$ (i.e., closer to iid), all methods perform better than $\alpha=0.1,$ as expected. With $\alpha=0.01,$ which is highly non-iid (c.f., Figure 5 in Appendix B), we see a significant drop in performance for all methods. SGD with LR decay and $\Delta$-SGD perform the best, while adaptive methods like Adam ($85\%\rightarrow 22.1\%$) and Adagrad ($84.1\%\rightarrow 12.5\%$) degrade noticeably more than other methods.
@@
For FMNIST trained with CNN (Figure 2(B)), which can be considered as an easier problem compared to CIFAR-10 classification, the performance degradation with varying $\alpha$’s is much milder. Interestingly, adaptive methods like Adam, Adagrad, and SPS perform much worse than other methods, indicating their inflexibility when applied to diverse datasets and model architectures, which often is worse than simple SGD-based methods; we further investigate this phenomenon in more detail in the next remark.
A similar trend can be seen in MNIST classification, where SGD-based methods work relatively well, except for SGDM with $\alpha=0.01$: without the LR decay, SGDM only achieves around 78% accuracy, while SGD without LR decay still achieves over $96\%$ accuracy (c.f., Figure 1(A)). Even with the LR decay, SGDM achieves noticeably worse performance. While this may seem like an “easy fix” in hindsight, it indicates that in a highly non-iid setting, one should be extremely careful in tuning SGD-based methods, even for easy datasets like MNIST. Interestingly, SPS performs particularly poorly, merely achieving 10% for MNIST with $\alpha=1,$ while achieving 96% for “harder” case of $\alpha=0.1,$ and again dropping to 69.7% with $\alpha=0.01$; a possible explanation is, with simple CNN, the model is not over-parameterized enough to have $f_{i}^{\star}=0,$ the default estimate we use on which the performance of SPS crucially rely Hazan and Kakade (2019). Adagrad similarly performs quite poorly.
@@
Lastly, in Figure 2(C), results for CIFAR-100 classification trained with ResNet-50 are illustrated. $\Delta$-SGD exhibits superior performance in all cases of $\alpha$. Unlike MNIST and FMNIST, Adam enjoys the second ($\alpha=0.1$) or the third ($\alpha=1$) best performance in this task, complicating how one should “tune” Adam for the task at hand. Other methods, including SGD with and without momentum/LR decay, Adagrad, and SPS, perform much worse than $\Delta$-SGD.
@@
Changing the model architecture. For this remark, let us focus on CIFAR-100 trained on ResNet-18 versus on ResNet-50, with $\alpha=0.1$, illustrated in Figure 3(A). SGD and SGDM (both with and without LR decay), Adagrad, and SPS perform worse using ResNet-50 than ResNet-18. This is a counter-intuitive behavior, as one would expect to get better accuracy by using a more powerful model. $\Delta$-SGD is an exception: without any additional tuning, $\Delta$-SGD can improve its performance. Adam also improves similarly, but the achieved accuracy is significantly worse than that of $\Delta$-SGD. While it is true that the performance of $\Delta$-SGD also slightly degrades for $\alpha=0.01,$ it still performs better than the other methods.
@@
Changing the dataset. We now focus on cases where the dataset changes, but the model architecture remains the same. We mainly consider two cases, illustrated in Figure 3(B) and (C): When CNN is trained for classifying MNIST versus FMNIST, and when ResNet-18 is trained for CIFAR-10 versus CIFAR-100. Aggravating the complexity of fine-tuning SGD(M) for MNIST, one can observe a similar trend in CIFAR-10 trained with ResNet-18. In that case, $\Delta$-SGD does not achieve the best test accuracy (although it is the second best with a pretty big margin with the rest), while SGD with LR decay does. However, without LR decay, the accuracy achieved by SGD drops more than 11%.
Interestingly, for the same setting, when we change the dataset from CIFAR-10 to CIFAR-100, SGD without LR decay (which is not one of the best methods for the CIFAR-10 case) achieves the second-best performance (after $\Delta$-SGD). On the other hand, SGDM with LR decay, which achieves the 3rd best performance for CIFAR-10 achieves very poor performance (more than 10% less than the best) for CIFAR-100. Finally, SGD with decay, which performs the best in the CIFAR-10 case, turns out to be the 3rd best algorithm when we change the dataset to CIFAR-100. Again, adaptive methods like Adam, Adagrad, and SPS perform quite poorly in this case.
@@
Figure 3: The effect of changing the dataset and the model architecture on different client optimizers. (A): CIFAR-100 trained with ResNet-18 versus Resnet-50 ($\alpha=0.1$), (B): MNIST versus FMNIST trained with CNN ($\alpha=0.01$), (C): CIFAR-10 versus CIFAR-100 trained with ResNet-18 ($\alpha=0.01$).
@@ 5 Conclusion
In this work, we propose $\Delta$-SGD, a distributed SGD scheme equipped with an adaptive step size that enables each client to use its step size and adapts to the local smoothness of the function each client is optimizing. We prove the convergence of $\Delta$-SGD in the general nonconvex setting and present extensive empirical results, where the superiority of $\Delta$-SGD is shown in various scenarios without any tuning. For future works, extending $\Delta$-SGD to a coordinate-wise step size in spirit of Duchi et al. (2011); Kingma and Ba (2014), as well as enabling asynchronous updates Assran et al. (2020); Toghani et al. (2022); Nguyen et al. (2022) could be interesting directions.
Broader Impacts and Limitations. We believe our proposed method, $\Delta$-SGD, can potentially impact reducing the amount of tuning required in training FL models. By now, it is widely believed that FL systems do not “dodge the bullet” of high carbon emission in AI/ML: as noted in a recent study Yousefpour et al. (2023), “compute on client devices, and the communication between the clients and the server are responsible for the majority of FL’s overall carbon emissions (97%).” The lack of principled hyperparameter tuning only exacerbates this problem. Based on the results presented in Section 4, $\Delta$-SGD achieves superior performance while robust to different settings.
In terms of limitations, our assumption on Theorem 1, especially Assumption 1c, as well as the conditions on $\gamma$ and $\eta_{0}$ are quite strong, and does not truly capture the strength of $\Delta$-SGD. Moreover, due to non-optimal implementation, $\Delta$-SGD takes about 1.3×1.3$\times$1.3 × more wall-clock time compared to SGD, which can be improved with more efficient implementation.
@@
    • In Section A.2, we provide the proof of Algorithm, 1, with some additional assumptions, including that $f_{i}$ is convex for all $i\in[m]$.
    • In Section A.3, we provide a simple proof in the centralized nonconvex setting, which was an open question posed in Malitsky and Mishchenko (2020).
@@ Appendix A Missing Proofs
We first introduce the following two inequalities which will be used in the following proofs.
@@
for any set of $m$ vectors $\{x_{i}\}_{i{=}1}^{m}$ such that $x_{i}\in\mathbb{R}^{d}$.
We also provide the following lemma, which establishes $\tilde{L}$-smoothness based on local smoothness, without assuming the global $L$-smoothness.
@@
For locally smooth function $f_{i}$, the following holds for the points that Algorithm 1 visits:
$f_{i}(x_{t})\leqslant f_{i}(x_{t-1})+\langle\nabla f_{i}(x_{t-1}),x_{t}-x_{t-1}\rangle+\frac{1}{4\eta_{t}^{i}}\|x_{t}-x_{t-1}\|^{2}.$ (9)
@@
By the definition of $\eta_{t}^{i},$ we have
$\|\nabla f_{i}(x_{t})-\nabla f_{i}(x_{t-1})\|\leqslant\frac{1}{2\eta_{t}^{i}}\|x_{t}-x_{t-1}\|$
@@
$\langle\nabla f_{i}(x_{t})-\nabla f_{i}(x_{t-1}),x_{t}-x_{t-1}\rangle$ $\leqslant\|\nabla f_{i}(x_{t})-\nabla f_{i}(x_{t-1})\|\cdot\|x_{t}-x_{t-1}\|$
$\leqslant\frac{1}{2\eta_{t}^{i}}\|x_{t}-x_{t-1}\|^{2}.$
Let $g(k)=f_{i}(x_{t-1}+k(x_{t}-x_{t-1})).$ Then, we have
$g^{\prime}(k)$ $=\langle\nabla f_{i}(x_{t-1}+k(x_{t}-x_{t-1})),x_{t}-x_{t-1}\rangle$
$g^{\prime}(k)-g^{\prime}(0)$ $=\langle\nabla f_{i}(x_{t-1}+k(x_{t}-x_{t-1})-\nabla f_{i}(x_{t}),\frac{k}{k}(x_{t}-x{t-1})\rangle$
$\leqslant\frac{k}{2\eta_{t}^{i}}\|x_{t}-x_{t-1}\|^{2}.$
@@
$\leqslant g(0)+g^{\prime}(0)+\frac{1}{4\eta_{t}^{i}}\|x_{t}-x_{t-1}\|^{2}$
$\leqslant f_{i}(x_{t-1})+\langle\nabla f_{i}(x_{t-1}),x_{t}-x_{t-1}\rangle+\frac{1}{4\eta_{t}^{i}}\|x_{t}-x_{t-1}\|^{2}.$
We also establish the averaged version of local $\tilde{L}$-smoothness. First, we re-write the statement of Lemma 2 as follows:
$\|\nabla f_{i}(x)-\nabla f_{i}(y)\|^{2}\leqslant\tilde{L}_{i}\|x-y\|^{2},$ (10)
where $\tilde{L}_{i}$ is $\max_{t}\Big{(}\frac{1}{2\eta_{t}^{i}}\Big{)}$ for all $x,y\in\text{conv}\{x_{0},x_{1},\dots\}$, and $\text{conv}\{\cdots\}$ denotes the convex hull of the arguments Boyd et al. (2004). Now, given (10),
@@
The average of $\tilde{L}$-local smooth functions $f_{i}$ are at least $\tilde{L}$-local smooth, where $\tilde{L}=\frac{1}{m}\sum_{i=1}^{m}\tilde{L}_{i}$.
@@
$\leqslant\frac{1}{m}\sum_{i=1}^{m}\tilde{L}_{i}\|x-y\|$
$=\tilde{L}\|x-y\|.$ (11)

### v3

@@ 4.1 Results
Table 1: Experimental results based on the settings detailed in Section 4. Best accuracy ($\pm$ 0.5%) for each task are shown in bold. Subscripts↓(x.x) is the performance difference from the best result, and is highlighted in pink when it is bigger than 2%. The down-arrow symbol ($\downarrow$) indicates step-wise learning rate decay, where the step sizes are divided by $10$ after 50%, and another by $10$ after 75% of the total rounds.
Figure 2: The effect of stronger heterogeneity on different client optimizers, induced by the Dirichlet concentration parameter $\alpha\in\{0.01,0.1,1\}$. $\Delta$-SGD remains robust performance in all cases, whereas other methods show significant performance degradation when changing the level of heterogeneity $\alpha$, or when changing the setting (model/architecture).
@@
Three illustrative cases are visualized in Figure 2. We remind that the step sizes are tuned for the task in Figure 2(A). For this task, with $\alpha=1$ (i.e., closer to iid), all methods perform better than $\alpha=0.1,$ as expected. With $\alpha=0.01,$ which is highly non-iid (c.f., Figure 8 in Appendix B.9), we see a significant drop in performance for all methods. SGD with LR decay and $\Delta$-SGD perform the best, while adaptive methods like Adam ($85\%\rightarrow 22.1\%$) and Adagrad ($84.1\%\rightarrow 12.5\%$) degrade noticeably more than other methods.
@@
Figure 2(B) shows the results of training FMNIST with a CNN. This problem is generally considered easier than CIFAR-10 classification. The experiments show that the performance degradation with varying $\alpha$’s is much milder in this case. Interestingly, adaptive methods such as Adam, Adagrad, and SPS perform much worse than other methods
@@
Lastly, in Figure 2(C), results for CIFAR-100 classification trained with a ResNet-50 are illustrated. $\Delta$-SGD exhibits superior performance in all cases of $\alpha$. Unlike MNIST and FMNIST, Adam enjoys the second ($\alpha=0.1$) or the third ($\alpha=1$) best performance in this task, complicating how one should “tune” Adam for the task at hand. Other methods, including SGD with and without momentum/LR decay, Adagrad, and SPS, perform much worse than $\Delta$-SGD.
@@
Changing the model architecture/dataset. For this remark, let us focus on CIFAR-100 trained on ResNet-18 versus on ResNet-50, with $\alpha=0.1$, illustrated in Figure 3(A). SGD and SGDM (both with and without LR decay), Adagrad, and SPS perform worse using ResNet-50 than ResNet-18. This is a counter-intuitive behavior, as one would expect to get better accuracy by using a more powerful model, although the step sizes are tuned on ResNet-18. $\Delta$-SGD is an exception: without any additional tuning, $\Delta$-SGD can improve its performance. Adam also improves similarly, but the achieved accuracy is significantly worse than that of $\Delta$-SGD.
We now focus on cases where the dataset changes, but the model architecture remains the same. We mainly consider two cases, illustrated in Figure 3(B) and (C): When CNN is trained for classifying MNIST versus FMNIST, and when ResNet-18 is trained for CIFAR-10 versus CIFAR-100. From (B), one can infer the difficulty in tuning SGDM: without the LR decay, SGDM only achieves around 78% accuracy for MNIST classification, while SGD without LR decay still achieves over $96\%$ accuracy. For FMNIST on the other hand, SGDM performs relatively well, but adaptive methods like Adam, Adagrad, and SPS degrade significantly.
@@
Aggravating the complexity of fine-tuning SGD(M) for MNIST, one can observe from (C) a similar trend in CIFAR-10 trained with ResNet-18. In that case, $\Delta$-SGD does not achieve the best test accuracy (although it is the second best with a pretty big margin with the rest), while SGD with LR decay does. However, without LR decay, the accuracy achieved by SGD drops more than 11%. Again, adaptive methods like Adam, Adagrad, and SPS perform suboptimally in this case.
@@
Figure 3: The effect of changing the dataset and the model architecture on different client optimizers. $\Delta$-SGD remains superior performance without additional tuning when model or dataset changes, whereas other methods often degrades in performance. (A): CIFAR-100 trained with ResNet-18 versus Resnet-50 ($\alpha=0.1$), (B): MNIST versus FMNIST trained with CNN ($\alpha=0.01$), (C): CIFAR-10 versus CIFAR-100 trained with ResNet-18 ($\alpha=0.01$).
Changing the domain: text classification. In Table 4(a), we make a bigger leap, and test the performance of different client optimizers for text classification tasks. We use the DBpedia and AGnews datasets (Zhang et al., 2015), and train a DistillBERT (Sanh et al., 2019) for classification. Again, the best step size for each client optimizer was tuned for the CIFAR-10 image classification trained with a ResNet-18, for $\alpha=0.1$. Thus, the type of the dataset changed completely from image to text, as well as the model architecture: from vision to language.
Not surprisingly, four methods (SGDM, SGDM($\downarrow$), Adam, and Adagrad) ended up learning nothing (i.e., the achieved accuracy is 1/(# of labels)), indicating the fine-tuned step sizes for these methods using CIFAR-10 classification task was too large for the considered text classification tasks. $\Delta$-SGD, on the other hand, still achieves competitive accuracies without additional tuning, thanks to the locality adaptive step size. Interestingly, SGD, SGD($\downarrow$), and SPS worked well for the text classification tasks, which is contrary to their suboptimal performances for image classification tasks in Table 1.
Changing the loss function. In Table 4(b), we illustrate how $\Delta$-SGD performs when combined with a different loss function, such as FedProx (Li et al., 2020). FedProx aims to address heterogeneity; we thus used the lowest concentration parameter, $\alpha=0.01$. We only present a subset of results: CIFAR-10(100) classification with ResNet-18(50) (additional results in Appendix B.6). The results show that $\Delta$-SGD again outperforms all other methods with a significant margin: $\Delta$-SGD is not only the best, but also the performance difference is at least 2.5%, and as large as 26.9%. Yet, compared to the original setting without the proximal term in Table 1, it is unclear whether or not the additional proximal term helps. For $\Delta$-SGD, the accuracy increased for CIFAR-10 with ResNet-18, but slightly got worse for CIFAR-100 with ResNet-50; a similar story applies to other methods.
Text classification Dataset / Model
$\alpha=1$ Agnews Dbpedia
Optimizer DistillBERT
SGD 91.1↓(0.5) 96.0↓(2.9)
SGD ($\downarrow$)↓ ) 91.6↓(0.0) 98.7↓(0.2)
SGDM 25.0↓(66.6) 7.1↓(91.8)
SGDM ($\downarrow$)↓ ) 25.0↓(66.6) 7.1↓(91.8)
Adam 25.0↓(66.6) 7.1↓(91.8)
Adagrad 25.0↓(66.6) 7.1↓(91.8)
SPS 91.5↓(0.1) 98.9↓(0.0)
$\Delta$-SGD 90.7↓(0.9) 98.6↓(0.3)
FedProx Dataset / Model
$\alpha=0.01$ CIFAR-10 CIFAR-100
Optimizer ResNet-18 ResNet-50
SGD 20.0↓(13.8) 25.2↓(5.9)
SGD ($\downarrow$)↓ ) 31.3↓(2.5) 20.2↓(10.8)
SGDM 29.3↓(4.4) 23.8↓(7.2)
SGDM ($\downarrow$)↓ ) 25.3↓(8.5) 15.0↓(16.0)
Adam 28.1↓(5.7) 22.6↓(8.4)
Adagrad 19.3↓(14.5) 4.1↓(26.9)
SPS 27.6↓(6.2) 16.5↓(14.5)
$\Delta$-SGD 33.8↓(0.0) 31.0↓(0.0)
Table 2: Additional experiments for a different domain (a), and a different loss function (b).
Other findings. We present all the omitted theorems and proofs in Appendix A, and additional empirical findings in Appendix B. Specifically, we demonstrate: the robustness of the performance of $\Delta$-SGD (Appendix B.1), the effect of the different numbers of local iterations (Appendix B.2), additional experiments when the size of local dataset is different per client (Appendix B.3), additional experimental results using FedAdam (Reddi et al., 2021) (Appendix B.4), additional experiential results using MOON (Appendix B.5) and FedProx (Appendix B.6) loss functions, and three independent trials for a subset of tasks (Appendix B.7). We also visualize the step size conditions for $\Delta$-SGD (Appendix B.8) as well as the degree of heterogeneity (Appendix B.9).
@@ 5 Conclusion
In this work, we proposed $\Delta$-SGD, a distributed SGD scheme equipped with an adaptive step size that enables each client to use its own step size and adapts to the local smoothness of the function each client is optimizing. We proved the convergence of $\Delta$-SGD in the general nonconvex setting and presented extensive empirical results, where the superiority of $\Delta$-SGD is shown in various scenarios without any tuning. For future works, extending $\Delta$-SGD to a coordinate-wise step size in the spirit of Duchi et al. (2011); Kingma & Ba (2014), applying the step size to more complex algorithms like (local) factored gradient descent (Kim et al., 2022a), and enabling asynchronous updates (Assran et al., 2020; Toghani et al., 2022; Nguyen et al., 2022) could be interesting directions.
Acknowledgments
This work is supported by NSF FET: Small No. 1907936, NSF MLWiNS CNS No. 2003137 (in collaboration with Intel), NSF CMMI No. 2037545, NSF CAREER award No. 2145629, NSF CIF No. 2008555, Rice InterDisciplinary Excellence Award (IDEA), NSF CCF No. 2211815, NSF DEB No. 2213568, and Lodieska Stockbridge Vaughn Fellowship.
@@ Acknowledgments
    • In Section A.2, we provide the proof of Algorithm, 1, with some modified assumptions, including that $f_{i}$ is convex for all $i\in[m]$.
@@
    • In Section B.4, we provide additional experimental results using FedAdam (Reddi et al., 2021) server-side adaptive method.
    • In Section B.5, we provide additional experimental results using MOON (Li et al., 2021) loss function.
    • In Section B.6, we provide additional experiemtns using FedProx (Li et al., 2020) loss function.
    • In Section B.7, we perform three independent trials for a subset of tasks for all client-optimizers. We plot the average and the standard deviations in Figure 6, and report those values in Table 7.
    • In Section B.8, we visualize the step size conditions for $\Delta$-SGD.
@@ Appendix A Missing Proofs
We first introduce the following two inequalities, which will be used in the subsequent proofs:
@@
for any set of $m$ vectors $\{x_{i}\}_{i{=}1}^{m}$ with $x_{i}\in\mathbb{R}^{d}$.
We also provide the following lemma, which establishes $\tilde{L}$-smoothness based on the local smoothness.
@@
For locally smooth function $f_{i}$, the following inequality holds:
$f_{i}(y)\leqslant f_{i}(x)+\langle\nabla f_{i}(x),y-x\rangle+\frac{\tilde{L}_{i,t}}{2}\|y-x\|^{2}\quad\forall x,y\in\mathcal{S}_{i,t},$ (9)
where $\mathcal{S}_{i,t}:=\mathrm{conv}(\{x_{t,k}^{i}\}_{k=1}^{K}\cup\{x_{t},x_{t+1}\})$.
@@
First, notice that by Assumptions (1a) and (1b), stochastic gradients are also bounded for all $x\in\mathbb{R}^{d}.$ Therefore, the local iterates, i.e., $\{x_{t,k}^{i}\}_{k=0}^{K}$ remain bounded. Hence, the set $\mathcal{S}_{i,t}:=\mathrm{conv}(\{x_{t,k}^{i}\}_{k=1}^{K}\cup\{x_{t},x_{t+1}\})$, where $\mathrm{conv}(\cdot)$ denotes the convex hull, is bounded. Due to the fact that $f_{i}$ is locally smooth, $\nabla f$ is Lipschitz continuous on bounded sets. This implies there exists a positive constant $\tilde{L}_{i,t}$ such that,
$\|\nabla f_{i}(x)-\nabla f_{i}(y)\|\leqslant\tilde{L}_{i,t}\|y-x\|\quad\forall x,y\in\mathcal{S}_{i,t}.$ (10)
@@
$\langle\nabla f_{i}(y)-\nabla f_{i}(x),y-x\rangle$ $\leqslant\|\nabla f_{i}(y)-\nabla f_{i}(x)\|\cdot\|y-x\|$
$\leqslant\tilde{L}_{i,t}\|y-x\|^{2},\quad\forall x,y\in\mathcal{S}_{i,t}.$
Let $g(\tau)=f_{i}(x+\tau(y-x)).$ Then, we have
$g^{\prime}(\tau)$ $=\langle\nabla f_{i}(x+\tau(y-x)),y-x\rangle$
$g^{\prime}(\tau)-g^{\prime}(0)$ $=\langle\nabla f_{i}(x+\tau(y-x))-\nabla f_{i}(x),y-x\rangle$
$\leqslant\|\nabla f_{i}(x+\tau(y-x))-\nabla f_{i}(x)\|\cdot\|y-x\|$
$\leqslant\tau\tilde{L}_{i,t}\|y-x\|^{2}.$
@@
$\leqslant g(0)+g^{\prime}(0)+\frac{\tilde{L}_{i,t}}{2}\|y-x\|^{2}$
$\leqslant f_{i}(x)+\langle\nabla f_{i}(x),y-x\rangle+\frac{\tilde{L}_{i,t}}{2}\|y-x\|^{2}.$
Now, we define $\tilde{L}_{i}:=\max_{t}\tilde{L}_{i,t}$. Given this definition, we relate the local-smoothness constant $\tilde{L}$ of the average function $f$ with (individual) local-smoothness constants $\tilde{L}_{i}$ in the below lemma.
@@
The average of $\tilde{L}_{i}$-local smooth functions $f_{i}$ are at least $(\frac{1}{m}\sum_{i=1}^{m}\tilde{L}_{i})$-local smooth.
@@
$\leqslant\frac{1}{m}\sum_{i=1}^{m}\tilde{L}_{i}\|x-y\|$ (11)
$\leqslant\tilde{L}\|x-y\|,\quad\text{where}\quad\tilde{L}=\max_{i,t}\tilde{L}_{i,t}.$ (12)
Technically, (11) can provide a slightly tighter bound for Theorem 1, but (12) simplifies the final expression. We also have the following form:
$f(x_{t+1})\leqslant f(x_{t})+\langle\nabla f(x_{t}),x_{t+1}-x_{t}\rangle+\frac{\tilde{L}}{2}\|x_{t+1}-x_{t}\|^{2},$
which can be obtained from Lemma 2. ∎